Given (triangle)ABC~(tringle)XYZ, AB=9, BC=18 and AC-21. If the perimeter of (triangle)XYZ is 112, find the measures of the three sides of (triangle)XYZ.
Answer by RAY100(1637)

Draw a rough sketch of two similar triangles
label sides carefully
as triangles are similar, sides are proportional
per of given triangle abc = 9 + 18 + 21 = 48.
per of new triangle is 112
as perimeters vary as scale factor to 1st power
sf = 112/48 =2 1/3 =7/3
xy = (7/3)ab =7/3 * 9 = 21
yz = (7/3) bc = 7/3*18 = 42
xz = (7/3) ac = 7/3*21 = 49
checking per 2nd triangle = 21+42+49= 112 ok